OpenStudy (anonymous):

Solve. 5x2 – 20 = 0 x = 5 and x = 4 x = 4 and x = 2 x = –5 and x = 20 x = 2 and x = –2

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

You can start by: Dividing both the sides by 5..

OpenStudy (anonymous):

5x^2-4=0

OpenStudy (raden):

x^2 - 4 = 0

OpenStudy (raden):

factorize by using the difference squares formula a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

And you can use here : \[x^2 - a^2 = (x + a) (x - a)\]

OpenStudy (anonymous):

Or if you don't wanna do this, then add 4 both the sides and then take square root both the sides.

OpenStudy (anonymous):

so,... whats the answer ?

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

its not B

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Who told you that answer is B?

OpenStudy (anonymous):

\(x^2 - 4 = 0\) Add 4 both the sides and tell me what did you get?

OpenStudy (raden):

well, look at how the factorization works x^2 - 4 = 0 x^2 - 2^2 = 0 use the diff square formula, (x+2)(x-2) = 0 solve for x find the zeroes by make each factor be zero first, x+2=0 subtract by 2 from both sides, x+2-2 = 0-2 x = -2 do for the rest!

OpenStudy (anonymous):

I dont really have to do the rest because there isnt another -2 in the other choices.

OpenStudy (raden):

hehe, good tricky :)

OpenStudy (anonymous):

do for the rest is actually for your own knowledge and not for answer choices you are given with..

OpenStudy (raden):

ok, show to us how to solve of x - 2 = 0 ?